OpenStudy (anonymous):

15. [3.05] Given the function g(x) = 8x – 2, find g(−2). (5 points) 14 0.5 –18 –14 16. [3.06] If f(x) = 2x2 − 30, find f(4). (5 points) 35 –14 2 17 17. [3.06] If h(x) = –4x − 10, find h(−5). (5 points) –30 1.25 10 –1 18. [3.06] Given the function f(x) = 2(x + 8), find x if f(x) = 12. (5 points) 2 40 –2 14

OpenStudy (anonymous):

Substitute the value of x in the equation

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

g(x) = 8x – 2 g(−2) = 8 *-2 -2 = -16-2 = -18

OpenStudy (anonymous):

Okay thank you @Yahoo!

OpenStudy (anonymous):

Other question help??

OpenStudy (anonymous):

I need help on the other answers........

OpenStudy (anonymous):

PLEASE?

OpenStudy (anonymous):

Same....Thing apply to other qquestion

OpenStudy (anonymous):

I do not understand @Yahoo!